Please answer with workDetermine all critical points for the function. 9) f(x) = x3 - 9x2 + 8 9) A) X = -3 and x = 3 B) x = 0 and x = 3 C) x = 0 D)
Answer #1

f\left ( x \right )=x^{3}-9x^{2}+8

f{}'\left ( x \right )=\frac{\mathrm{d} }{\mathrm{d} x}\left [x^{3}-9x^{2}+8 \right ]

=\frac{\mathrm{d} }{\mathrm{d} x}\left [x^{3} \right ]-\frac{\mathrm{d} }{\mathrm{d} x}\left [9x^{2} \right ]+\frac{\mathrm{d} }{\mathrm{d} x}\left [8 \right ]

=3x^{2} -9\times \frac{\mathrm{d} }{\mathrm{d} x}\left [x^{2} \right ]+0

=3x^{2} -9\times 2x

=3x^{2} -18x

f{}'\left ( x \right )=0

\Rightarrow 3x^{2} -18x=0

\Rightarrow 3x\left ( x-6 \right )=0

x\left ( x-6 \right )=0

x=0 or x-6=0

x=0 or x=6

That is

Function have critical points at x=0,6

Correct option is

------------------------------------------------------------------